OpenStudy (anonymous):

Find dy/dx: ye^x-x=y^2

myininaya (myininaya):

y'e^x+ye^x-1=2yy' y'e^x-2yy'=1-ye^x y'(e^x-2y)=1-ye^x y'=(1-ye^x)/(e^x-2y)